40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Sun Sep 9 01:18:23 2012, tara, Notes, Schedule, Task Sheet tasksheet.pngtask_sheet_2.xlsx
    Reply  Thu Jan 24 16:21:01 2013, tara, Notes, Schedule, Task Sheet Thermal_Noise_Probe_Schedule_-_Sheet1-2.pdf
       Reply  Mon Feb 18 03:38:20 2013, tara, Notes, Schedule, Task Sheet 
Message ID: 1101     Entry time: Mon Feb 18 03:38:20 2013     In reply to: 1094
Author: tara 
Type: Notes 
Category: Schedule 
Subject: Task Sheet 

Now that the layout is ok and the table is cleared, the next thing to do is open the chamber. The plan for this week will be:

==Cavities & Vac chamber==

  1. clear the clean bench
  2. Prepare all parts  for the new cavities that will go to the chamber. The assembly for the new cavities is almost done, I'm wating for the caps of the thermal shield.
  3. Open the chamber, remove the 8" cavities.
  4. Change the heaters/ thermometers. Currently, the feedthrough for the chamber is 9-pin. It is for 3 thermometer and 1 heater. The new setup will need 2 thermometers, 2 heaters. I don't know what kind of wire and where it is in the lab yet. This need to be checked.
  5. Once the sensors/heaters are installed, we can close the chamber and start the mode matching.

==1st laser==

  1.  The setup for the first laser should be easy. The setup from the laser to the PMC is not changed much except the BB eom for locking. I should be able to lock the first cavity soon

==2nd laser==

  1. Need to think about the PMC
  2. The EOM
ELOG V3.1.3-